OpenStudy (gabylovesyou):

Admission to the zoo costs $4 for adults and $2 for kids. The total bill for your entire extended family was $30. This scenario can be modeled by the equation 4a + 2k = 30. Which is not a possible combination of adults and kids? 5 adults and 5 kids 4 adults and 7 kids 7 adults and 1 kid 3 adults and 6 kids

OpenStudy (shane_b):

For this one you could simply try the values and see which equation is wrong:\[4a+2k=30\]\[4(5)+2(5)=30\]\[4(4)+2(7)=30\]\[4(7)+2(1)=30\]\[4(3)+2(6)\neq30\]

OpenStudy (gabylovesyou):

so its D

OpenStudy (shane_b):

Yes

OpenStudy (gabylovesyou):

sank u very much

OpenStudy (shane_b):

yw